A pinball machine has five holes of varying sizes in which the ball may land. The payoff for the


ball landing in a given hole indicates that the probabilities of success for various holes should be


0.28, 0.24, 0.19, 0.15 and 0.14.


A contestant plays the game 42 times and obtains the following frequencies: 15, 10, 7, 6 and 4. Test


at a 5% level of significance whether the probabilities corresponding to the payoffs are correct.

1
Expert's answer
2021-11-10T15:38:18-0500

This question requires us to perform a goodness of fit test. To perform this test, we shall use the Chi-square distribution for goodness of fit.

The hypotheses tested are,

"H_0:"  probabilities corresponding to the payoffs are correct.

"Against"

"H_1:"  probabilities corresponding to the payoffs are not correct.

The expected frequencies "(E_i)" are first determined using the formula,

"E_i=n*p_i" where "n=15+10+7+6+4=42" and "p_i" are the probabilities of success for each observed count "O_i" .

The expected frequencies are given as,

"E_1=n*p_1=42*0.28=11.76"

"E_2=n*p_2=42*0.24=10.08"

"E_3=n*p_3=42*0.19=7.98"

"E_4=n*p_4=42*0.15=6.3"

"E_5=n*p_5=42*0.14=5.88"

A summary of the above information is given in the table below

"O_i" "p_i" "E_i"

15 0.28 11.76

10 0.24 10.08

7 0.19 7.98

6 0.15 6.3

4 0.14 5.88

Where "i=1,2,3,4,5" is the number of holes.

The test statistic is given as,

"\chi^2=\displaystyle\sum^5_{i=1}(O_i-E_i)^2/E_i"

"\chi^2=(15-11.76)^2/11.76+(10-10.08)^2/10.08+(7-7.98)^2/7.98+(6-6.3)^2/6.3+(4-5.88)^2/5.88= 1.629013"

"\chi^2" is compared with the table value at "\alpha=5\%" with "(i-1)=5-1=4". The table value is "\chi^2_{\alpha,4}=\chi^2_{0.05,4}=9.48773" and the null hypothesis is rejected if "\chi^2\gt \chi^2_{0.05,4}"

Since "\chi^2=1.629013\lt \chi^2_{0.05,4}=9.48773", we fail to reject the null hypothesis and we conclude that sufficient evidence exist to support the claim that probabilities corresponding to the payoffs are correct at 5% level of significance.